OpenStudy (anonymous):

solve. -15(4 + 3x) - 8x = 26 - 6 (x - 9)

OpenStudy (cwrw238):

first distribute -15 over the (4 + 3x) and -6 over (x-9) can you do that?

OpenStudy (anonymous):

swo it would be -15/7x and -15x

OpenStudy (cwrw238):

no - you multiply 4 by -15 then +3 by -15 distributive law is a(b + c) = ab + ac

OpenStudy (anonymous):

-60 and -45

OpenStudy (cwrw238):

so it -15(4 + 3x) = -60 - 45x

OpenStudy (cwrw238):

sorry - left out the x after the 3

OpenStudy (cwrw238):

so the equation becomes -60 - 45x - 8x = 26 - 6x + 54

OpenStudy (cwrw238):

do u follow that ok?

OpenStudy (anonymous):

a little

OpenStudy (cwrw238):

right - generally you work out the parentheses first

OpenStudy (anonymous):

ok

OpenStudy (cwrw238):

-6(x - 9) = -6*x -6*-9 = -6x + 54 (note : - times - = +)

OpenStudy (cwrw238):

writing it out again -60 - 45x - 8x = 26 - 6x + 54 now we need to bring all terms in x (like -6x) to the left hand side of the = and all the nmnbers to right hand side so add 60 to both sides -60 - 45x - 8x + 60 = 26 - 6x + 54 + 60 60's on left hand side cancel out -60 + 60 = 0 so; - 45x - 8x = 26 - 6x + 54 + 60 now add -6x to both sides ( to get rid of -6 on RHS)

OpenStudy (cwrw238):

* typo add 6x

OpenStudy (cwrw238):

any questions - ru following this ok?

OpenStudy (anonymous):

im following so far

OpenStudy (cwrw238):

good adding 6x to both sides: copying from last post - 45x - 8x = 26 - 6x + 54 + 60 -45x - 8x + 6x = 26 -6x + 6x + 54 + 60 6x's cancel out on RHS so we get - 45x - 8x + 6x = 26 + 54 + 60 adding -43x = 140

OpenStudy (anonymous):

that's a little confusing but i got it thank you so much for helping me

OpenStudy (cwrw238):

yes - practicing these will help you grasp the ideas behind the algebra i made a slight mistake on last line its 47x = 140 not 43x and x is found by dividing both sides by -47 so x = 140 / -47 or -140/47

OpenStudy (cwrw238):

no probs monica

OpenStudy (cwrw238):

to practice try an easier one like 4(x + 1) = 8

OpenStudy (cwrw238):

if you want to!!1

OpenStudy (anonymous):

so to find the answer i have to divide 140/-7

OpenStudy (cwrw238):

yes 140/-47 use calculator

OpenStudy (cwrw238):

you should get - 2.98 to decimal places

OpenStudy (anonymous):

thats what i got

OpenStudy (cwrw238):

good